### v2

@@ Abstract
Federated graph learning is an emerging field with significant practical challenges. While many algorithms have been proposed to enhance the accuracy of training graph neural networks, e.g., for node classification problems on large graphs, in a federated manner, their system performance is often overlooked, even though it is crucial for real-world deployment. To address this gap, we introduce FedGraph, a research library built for practical distributed deployment and benchmarking in federated graph learning. FedGraph supports a range of state-of-the-art graph learning methods and includes built-in profiling tools to evaluate system performance, focusing specifically on communication and computation costs during training. Unlike existing benchmark platforms, FedGraph natively incorporates homomorphic encryption to enhance privacy preservation and facilitates the development of practical applications by enabling distributed training across multiple physical machines, providing an evaluation framework that can guide the system design of future federated graph learning algorithms. Leveraging these optimizations, we use FedGraph to demonstrate the first privacy-preserving federated learning system to run on graphs with 100 million nodes.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1 Introduction
Graph neural networks aim to learn representations of graph-structured data that capture features associated with graph nodes as well as edges between the nodes Bronstein et al. (2017). Such representations can be used to represent and classify nodes or predict the formation of links in large graphs, or can be used to represent and classify smaller graphs. Popular applications of these learning tasks range from recommendation systems (e.g., predicting link formation in graphs if links represent user preferences for products) and fake news detection in social networks (e.g., classifying nodes in a social network graph as “fake” where nodes represent news items) to drug discovery in biology (e.g., representing cells as graphs and classifying them according to drug susceptibility) and fraud detection in sensor networks (e.g., classifying nodes as “fraudulent” if the nodes represent sensors connected together in a graph) Benamira et al. (2019); Zhang et al. (2020b).
@@
Real-world graph data is often too large to be trained on a single server or may naturally exist on multiple local clients. For example, graph learning on records of billions of users’ website visits requires significant computational resources beyond those of a single server. Even if a single server or data center could hold such information, privacy regulations may require that it be stored in the computers or smartphones at which it was generated. Strict data protection regulations such as the General Data Protection Regulation (GDPR) in Europe and Payment Aggregators and Payment Gateways (PAPG) in India, for example, prevent private user data from being shared across international borders. Users may also not want to share their personal data with the server. Federated learning then shows promise for preserving user privacy while training accurate models on data stored at multiple clients Zhao et al. (2018).
@@
These challenges to training on real-world graphs motivate federated graph learning (FGL) as an important research topic Yao et al. (2024); Xie et al. (2021); Zhang et al. (2021). The training schematic of such a federated architecture is shown in Figure 1: each client maintains a local graph and a local graph neural network (GNN) model, which can be aggregated with other clients’ local models at a coordinator server. Within this architecture, there are many different FGL algorithms that handle local model updates and aggregation differently, resulting in varying model accuracies and system performance (e.g., runtime, communication costs) Yao et al. (2024); Xie et al. (2021); Zhang et al. (2021). While evaluating the model accuracy of such algorithms can be easily done with simulations on open graph datasets Hu et al. (2020), evaluating their real-world system performance requires sophisticated benchmarking platforms that can fairly compare multiple FGL algorithms.
@@
Vanilla FGL benchmark platforms are designed for research purposes only and do not have a distributed training system, instead of simulating the actions of multiple physically distinct clients on a single machine. Thus, they do not capture the real communication and computation costs of training graph models in a federated manner. Meanwhile, existing federated learning libraries mainly focus on overcoming client data heterogeneity with limited support for GNN training. For example, FedScale Lai et al. (2022) does not support graph models, while FedGraphNN He et al. (2021) and FederatedScope-GNN Wang et al. (2022) are not well maintained and have limited support of multiple FGL algorithms, forcing users to implement the FGL algorithms that they desire to compare on these benchmark platforms themselves. Moreover, none of these platforms natively support FGL enhancements that may be needed in practice, e.g., homomorphic encryption for enhanced privacy preservation Jin et al. (2023). Table 1 compares their features to our FedGraph library.
Figure 1: Federated Graph Training Schematic. Nodes in a graph (shown as circles) are distributed across clients. Arrows indicate that each client can exchange model updates with a central server during the training process of a graph neural network model.
Vanilla FGL FedGraphNN FedScale FederatedScope-GNN FedGraph
@@
System Level Profilor
Large Graph Node Classification
Table 1: FedGraph Comparison with Former Frameworks. FedGraph supports distributed training with various FGL algorithms. It then supports algorithms for overcoming cross-client edges and encrypted aggregation. With a system-level profiler, FedGraph can then benchmark existing algorithms and provide optimizations to support node classification on large graphs.
To meet these shortcomings, we introduce FedGraph (Federated Graph) [^1] https://github.com/fedgraph/fedgraph, a research library to easily train GNNs in federated settings. As shown in Table 1, FedGraph supports various federated training methods of graph neural networks under both simulated and real federated environments, as well as encrypted communication between clients and the central server for model update and information aggregation. We summarize our contributions in FedGraph as follows.
  • We introduce FedGraph, a comprehensive Python library built for real-world federated graph learning systems, incorporating multiple state-of-the-art FGL algorithms to facilitate easy comparison between them.
  • FedGraph provides an advanced system-level profiling tool that enables detailed analysis of communication and computation costs throughout training, as well as encrypted aggregation to strengthen privacy protection.
  • Experimental benchmarks highlight FedGraph’s effectiveness in addressing key challenges in federated graph learning. We demonstrate that FedGraph enables detailed comparisons of different FGL algorithms under various configurations and showcases its ability to support a privacy-preserving federated system on graphs with 100 million nodes.
We give an overview of FedGraph’s system design in Section 2 before detailing our implementation of FedGraph’s components in Section 3. We discuss the usability of FedGraph as an evaluation platform in Section 4 and present examples of the experiments that it enables in Section 5. We then conclude in Section 6.
@@ 2 FedGraph System Design
In this section, we first introduce the three main graph learning problems that FedGraph supports and highlight the design principles of the library and how we achieve them.
2.1 Graph Learning Problems
Most FGL applications can modeled as one of three major graph learning problems: node classification (e.g., classifying nodes representing papers in citation networks based on their research topic), link prediction (e.g., recommending the formation of links that represent friendship between users in social networks), and graph classification (e.g. classifying types of proteins in biology, where each protein is represented as a graph). Figure 2 illustrates the three major aspects as different graph tasks.
Figure 2: Modeling Applications as Graph Tasks. Nodes are connected by edges, where the color of the node represents the information of the node. Node classification (left) is to classify the type of grey nodes based on the graph structure and the information from other nodes. Link prediction (middle) is to predict the future links between nodes based on existing edge and node information. Graph classification (right) is to classify the type of the graph based on the graph structure.
Given these three learning problems, FedGraph supports running algorithms that aim to solve the following learning tasks:
Federated Node Classification with Cross-Client Edges: A key challenge in node classification is the presence of cross-client edges that connect nodes at different clients Yao et al. (2024). Thus, the library supports communicating information about these nodes stored in other clients, either before or during training, without affecting the privacy of users.
Federated Link Prediction on Dynamic Graphs: A key challenge in link prediction is that links may form at clients at different rates. Thus, the library supports balancing temporal heterogeneity across clients with privacy preservation.
Federated Graph Classification on Heterogeneous Data: The library supports various algorithms for graph classification with non-IID distributions of graphs across clients.
@@ 2.2 Design Principles
The library is designed for a federated learning researcher or applied scientist, as shown in the design diagram of FedGraph in Figure 3. Our design follows four principles:
@@
Figure 3: Design Diagram of FedGraph. The system is organized into four layers: the user access layer, the application layer, the domain layer, and the infrastructure layer. Users only need to focus on the access layer while the developers can focus on one of the remaining layers based on the domain knowledge.
Easy-to-use and unified API: In the access layer, users first set up the training configuration and then use the same code for local simulation and federated training. All it takes is 10-20 lines of code to get started with training a federated GNN model with the same code for both simulation and distributed training. GNN models are PyTorch models provided by PyG and DGL. The federated training process is handled by Ray. We abstract away the complexity of federated graph training and provide a unified API for training and evaluating FedGraph models.
Various federated graph learning methods: In the application layer, the library supports three tasks with various algorithms and datasets. Most of the state-of-the-art federated graph training methods have been implemented by library developers or authors of research papers and are ready to be applied.
Great flexibility on new datasets and new algorithms: For the domain layer, we separate the code into data loader and federated training classes for better management. Existing FedGraph models can easily be extended for conducting your research. Simply inherit the base class of trainers and implement your methods.
Large-scale real-world federated graph training: In the infrastructure layer, by utilizing Ray and Kubernetes, the library can execute processes and store data in various edge/cloud computation devices. Homomorphic encrypted computation can be enabled to further enhance the federated aggregation. We can then meet the demand for FedGraph applications in challenging real-world scenarios with privacy preservation and support learning on large-scale graphs across multiple clients.
3 FedGraph System Components
In this section, we first present the FedGraph API for running different federated graph learning tasks. We then introduce FedGraph Homomorphic Encryption for privacy-preserved aggregation and FedGraph Kubernetes for scalable distributed training.
3.1 FedGraph API
Figure 4: High-level Demonstration of FedGraph API Design
@@ 3.2 FedGraph Homomorphic Encryption
FedGraph supports homomorphic encryption (HE) to ensure end-to-end secure computation to preserve user privacy across various graph learning algorithms, as shown in Figure 5. To our knowledge, this is the first federated graph learning library to incorporate end-to-end homomorphic encryption. While previous work explored homomorphic encryption in specific federated graph learning context Yao et al. (2024), FedGraph advances this approach through several key innovations that address challenges:
1. The system implements HE in different stages of pre-training and training based on algorithm design, enabling flexible security protocols across different FGL tasks.
2. The system maintains an efficient implementation of the homomorphic encryption process.
3. The system ensures secure aggregation for both feature and structural information in graph data
Figure 5: Homomorphic Encryption Process. The clients own the private key and the server is used to perform secure aggregation without knowing the plaintext data.
Pre-Train Aggregation: During pre-training, each client maintains its local node features and needs to share the aggregated neighbor information without revealing the raw data. Such information is encrypted before sending it to the server. The server then performs encrypted aggregation and then sends the result back to clients. The client then decrypts the feature aggregation without knowing the original feature.
Federated Aggregation During Training: During model aggregation on the server side, HE can also ensure secure aggregation during federated training without revealing fine-grained local updates to the server Zhang et al. (2020a); Jin et al. (2023). For the training process, we encrypt gradient updates before transmission for training parameters. Then, the central server performs aggregation on the encrypted values and homomorphic averaging on the aggregation. Such HE aggregation provides robust protection while maintaining computational efficiency.
@@ 3.3 FedGraph Kubernetes
We employ Kubernetes to optimize the scalability and reliability of the application layer, allowing our federated learning system to handle dynamic workloads effectively. As illustrated in Figure 6, our approach utilizes Amazon Web Services (AWS) Elastic Kubernetes Service (EKS) to facilitate flexible resource management, deploying a self-managed Kubernetes cluster equipped with GPU instances to support computationally intensive tasks. The architecture comprises a master node that oversees orchestration responsibilities, including task scheduling, resource allocation, and monitoring of cluster health, while worker nodes are responsible for executing distributed processes in parallel. To further enhance scalability and efficiency, we integrate Kubernetes Cluster Autoscaler, which automatically adjusts the number of worker nodes in response to fluctuating workload demands. This setup ensures that resources are provisioned only when necessary, leading to optimized costs and performance. The result is a robust, scalable federated learning environment that maintains high efficiency even under varying levels of load, supporting continuous, large-scale learning processes.
Figure 6: Kubernetes Service for FedGraph Computation Cluster Management
3.4 FedGraph Monitoring System
The FedGraph Monitoring System, represented in Figure 7, uses a Monitor Class to track CPU, GPU usage, memory, and communication costs between server and trainer nodes. Here is an overview of the monitoring system:
CPU/GPU Usage Monitoring: The Monitor Class keeps track of resource consumption within the Server Process and Ray Pod AutoScaler, providing insights into processing efficiency and identifying when additional resources might be required.
Memory Usage Tracking Memory usage across the API Gateway, Object Store, and server processes are monitored to detect potential bottlenecks, helping to ensure smooth operation and efficient memory allocation.
Communication Cost Logging: Data transfer rates between the FedGraph User and API Gateway are recorded to measure network performance and detect any latency issues affecting responsiveness.
This system optimizes resource use, tracks performance trends, and helps identify bottlenecks in real time.
Figure 7: FedGraph Monitoring System Architecture
3.5 Scalability
The FedGraph system is built to handle large-scale federated learning, supporting thousands to millions of nodes through the following key design features:
Client Selection for Large-Scale Federated Learning: FedGraph uses a selective client engagement mechanism, only partial clients send the model update at each training round. This approach reduces communication costs and resource strain, enabling scalability to thousands of clients per training round.
Minibatch Training for Federated Updates: Minibatch training allows each client to process only a subset of the local graph, reducing computational and memory demands. This enables devices with limited resources to participate, improving convergence speed without overwhelming network resources.
Communication and Resource Optimization: FedGraph optimizes communication, minimizing data transfer between clients and the server. Additionally, Kubernetes-based dynamically manages Server Pods and Trainer Pods to meet workload demands, ensuring efficient resource usage as the system scales.

### v3

@@ Abstract
Federated graph learning is an emerging field with significant practical challenges. While algorithms have been proposed to improve the accuracy of training graph neural networks, such as node classification on federated graphs, the system performance is often overlooked, despite it is crucial for real-world deployment. To bridge this gap, we introduce FedGraph, a research library designed for practical distributed training and comprehensive benchmarking of FGL algorithms. FedGraph supports a range of state-of-the-art graph learning methods and includes a monitoring class that evaluates system performance, with a particular focus on communication and computation costs during training. Unlike existing federated learning platforms, FedGraph natively integrates homomorphic encryption to enhance privacy preservation and supports scalable deployment across multiple physical machines with system-level performance evaluation to guide the system design of future algorithms. To enhance efficiency and privacy, we propose a low-rank communication scheme for algorithms like FedGCN that require pre-training communication, accelerating both the pre-training and training phases. Extensive experiments benchmark different FGL algorithms on three major graph learning tasks and demonstrate FedGraph as the first efficient FGL framework to support encrypted low-rank communication and scale to graphs with 100 million nodes [^1] Code available at https://github.com/FedGraph/fedgraph.
@@ 1 Introduction
Graph neural networks aim to learn representations of graph-structured data that capture features associated with graph nodes as well as edges between them (Bronstein et al., 2017). Such representations can be used to represent and classify nodes or predict the formation of links in large graphs, or can be used to represent and classify smaller graphs. Most graph applications can modeled as one of three major graph learning problems: node classification (e.g., classifying nodes representing papers in citation networks based on their research topic), link prediction (e.g., recommending the formation of links that represent friendship between users in social networks), or graph classification (e.g. classifying types of proteins in biology, where each protein is represented as a graph). Figure 1 (left) illustrates these three types of graph learning tasks (Benamira et al., 2019; Zhang et al., 2020b).
@@
In practice, graph data is often too large to be trained on a single server or may naturally exist on multiple local clients. For example, graph learning on records of billions of users’ website visits requires significant computational resources, beyond those of a single server. Even if a single server or data center could hold such information, privacy regulations may require that it be stored where it was generated, e.g., the General Data Protection Regulation (GDPR) in Europe and Payment Aggregators and Payment Gateways (PAPG) in India prevent private user data from being shared across international borders. Users may also not want to share their personal data with an external server. In response to these concerns, federated learning has been widely studied as a way to preserve user privacy while training accurate models on data stored at multiple clients (Zhao et al., 2018).
Figure 1: Modeling Applications as Graph Tasks (left) and Federated Graph Training Schematic (right). Node classification aims to predict the labels of grey nodes using structural and feature information from neighboring nodes. Link prediction involves inferring potential future edges between nodes based on existing graph structure and node features. Graph classification assigns labels to entire graphs based on their topology and node information. In federated graph learning, nodes are located across multiple clients with potential cross-client edges. Each client trains a local GNN on its subgraph and exchanges encrypted model updates with a central server, enabling collaborative learning without sharing raw data.
@@
These challenges to training on real-world graphs motivate Federated Graph Learning (FGL) as an important research topic (Liu et al., 2024). The training schematic of such a federated architecture is shown in Figure 1 (right): each client maintains a local graph and a local Graph Neural Network (GNN) model, which can be aggregated with other clients’ local models at a coordinator server. Within this architecture, there are many different FGL algorithms that handle local model updates and aggregation differently, resulting in varying model accuracies and system performance (e.g., runtime, communication costs) (Xie et al., 2021; Zhang et al., 2021; Yao et al., 2024). While evaluating the model accuracy of such algorithms can be easily done with simulations on open graph datasets (Hu et al., 2020), evaluating their real-world system performance requires sophisticated benchmarking platforms that can fairly compare multiple FGL algorithms.
@@
Vanilla FGL benchmark platforms are generally designed for research purposes only and simulate the actions of multiple clients on a single machine (Xie et al., 2021; Li et al., 2024). Thus, they do not capture the real communication and computation costs of training graph models in a federated manner across distinct physical clients. Meanwhile, existing federated learning libraries mainly focus on overcoming client data heterogeneity, with limited support for GNN training. For example, FedScale (Lai et al., 2022) does not support graph models, while FedGraphNN (He et al., 2021) and FederatedScope-GNN (Wang et al., 2022) are not well-maintained and only support basic federated learning methods like FedAvg. Users need to implement the FGL algorithms themselves, including steps unique to FGL tasks such as handling cross-client graph edges that begin and end at different clients. Moreover, none of these platforms natively support FGL enhancements that may be needed in practice, e.g., homomorphic encryption for enhanced privacy (Jin et al., 2023; Yao et al., 2024).
To meet these shortcomings, we introduce FedGraph, a research library to easily train GNNs in federated settings. As shown in Table 1, FedGraph supports various federated training methods of graph neural networks under both simulated and real federated environments, as well as encrypted communication between clients and the central server for model update and information aggregation. We summarize the contributions of FedGraph as follows.
  • FedGraph is the first Python library built for real-world federated graph learning systems, including system optimizations that improve efficiency, scalability, and privacy preservation, as well as multiple state-of-the-art FGL algorithms for easy comparison between them.
  • FedGraph natively supports homomorphic encrypted aggregation to strengthen privacy protection and provides an advanced system-level monitor class that enables detailed analysis of communication and computation costs throughout training.
  • FedGraph proposes a low-rank communication scheme for algorithms like FedGCN to accelerate both pre-train communication and training phases.
  • Extensive experiments benchmark the system performance of various algorithms on three major FGL tasks and showcase the ability to support a privacy-preserving federated system on graphs with 100 million nodes.
Vanilla FGL FedScale FedGraphNN FederatedScope-GNN FedGraph
@@
System Level Profiler
Large Scale ML Optimizations
Table 1: FedGraph Comparison with Existing Frameworks. FedGraph supports distributed training with various FGL algorithms, including algorithms that specifically support cross-client edges and encrypted aggregation. With a system-level profiler, FedGraph can then benchmark existing algorithms and provide optimizations to support training on large (tens of millions of nodes) graphs.
In this paper, we begin with an overview of the system design in Section 2, followed by highlighting the key system components in Section 3. In Section 4, we present a case study demonstrating how FedGraph facilitates the easy design and test of low-rank pre-training communication in FGL. We then benchmark the performance on three major tasks and evaluate its scalability in Sections 5. We finally conclude the paper in Section 6.
@@ 2 FedGraph System Design
In this section, we outline the design principles of the FedGraph library and demonstrate how these principles are implemented in our system design (Figure 2).
@@ 2.1 Design Principles
The main focus of FedGraph is providing a scalable and privacy-preserving federated graph learning system with ease of use for federated learning researchers and applied scientists in industry. As illustrated in Figure 2 (left), the system architecture is structured according to four design principles.
Optimized usability with simple configurations for simulation and federated training: At the access layer, users configure training settings and can seamlessly switch between local simulation and federated training using the same codebase. With just 10–20 lines of code, users can train a federated GNN model, as we discuss further in Section 2.2. FedGraph abstracts away the complexity of federated graph training, offering a unified platform for training and evaluation.
Benchmark existing federated graph learning methods: In the application layer, FedGraph supports three FGL tasks (node classification, link prediction, and graph classification) and a wide range of state-of-the-art model training algorithms for each task, implemented either by the original authors or library developers. Appendix C lists the currently supported algorithms and datasets.
Extensibility to new datasets and algorithms: In the domain layer, FedGraph separates code into modular components, such as data loaders and training classes. Researchers can extend existing models by inheriting base trainer classes and implementing their own algorithms or data pipelines.
Scalability and privacy preservation for real-world deployment: At the infrastructure layer, FedGraph leverages Ray and Kubernetes to enable distributed computation across edge and cloud environments. Homomorphic encryption can be enabled to secure federated aggregation. This architecture ensures that the system can meet the demands of privacy-preserving, large-scale federated graph learning in realistic settings.
@@
Figure 2: Design Diagram of FedGraph (left) and Quick Start Example (right). The system is organized into four layers: the user access layer, the application layer, the domain layer, and the infrastructure layer. Users only need to focus on the access layer, while the developers can focus on one of the remaining layers based on the domain knowledge.
2.2 FedGraph Use Example
Researchers can easily install the FedGraph library by running pip install fedgraph in Python, configure experiments, and start training federated GNN models. Appendices AB, and D provide more details about the code structure and user API. As illustrated in Figure 2 (right), users can initiate a federated graph learning experiment through a straightforward configuration process. Task, method, and dataset specification: The user specifies the learning task, algorithm, dataset, and client data distribution. FedGraph enforces explicit task-method combinations to ensure compatibility and reproducibility. Training configuration: Key hyperparameters, such as the number of global rounds, local training steps, learning rate, and number of participating trainers, are defined in the configuration. Security and privacy settings: Users can enable homomorphic encryption to ensure privacy-preserving training. Execution: Once settled configuration, calling the API function run_fedgraph(config) launches the experiment. FedGraph then automatically handles data loading, client initialization, and distributed training across clients.
3 FedGraph System Highlights
In this section, we first introduce FedGraph’s monitoring system, which enables usable benchmarking of FGL methods. We then introduce two infrastructure features, FedGraph Homomorphic Encryption for privacy-preserving aggregation and FedGraph Kubernetes for scalable distributed training. Finally, we discuss supported configurations for users to optimize large-scale model training.
3.1 FedGraph Monitoring System
Figure 3: FedGraph Monitoring System Architecture.
The FedGraph Monitoring System, illustrated in Figure 3 (right), uses a Monitor Class to track key system metrics, including running time, CPU/GPU utilization, memory consumption, and communication costs between the server and clients. Examples of the monitoring dashboard during training are provided in Appendix E.
Communication Cost Logging: The system records data transfer rates between the FedGraph User and the API Gateway to assess network performance and identify potential bottlenecks.
Training Time and Accuracy Logging: The monitoring framework captures the training duration and model accuracy throughout the learning process, enabling comparisons across configurations.
CPU/GPU and Memory Usage Tracking: Resource consumption across components, such as the Server Process, Ray Pod AutoScaler, API Gateway, and Object Store, is continuously monitored to provide insights into processing efficiency and to detect memory bottlenecks or potential leaks.
Section 5 further illustrates FedGraph’s resource usage profiles under various experimental settings.
Figure 4: FedGraph Homomorphic Encryption for Secure Aggregation on Encrypted Data.
@@ 3.2 FedGraph Homomorphic Encryption
FedGraph supports homomorphic encryption (HE) to enable end-to-end secure computation, ensuring user privacy during training (Figure 4). To our knowledge, it is the first FGL library to incorporate end-to-end HE as a core capability. While prior work has applied HE to specific FGL scenarios (Yao et al., 2024), FedGraph generalizes this support by addressing two key requirements of realistic FGL settings: (i) secure feature aggregation prior to training (Yao et al., 2024), and (ii) secure model aggregation during training (Zhang et al., 2024; Kim et al., 2025).
Pre-Training Aggregation: For algorithms such as FedGCN that require feature aggregation before training, FedGraph applies HE to protect data privacy. Each client encrypts its local node features and transmits the encrypted data to the server. The server performs aggregation directly on the ciphertext and returns the result, which clients then decrypt to obtain the aggregated neighbor features without knowing the original feature.
Federated Aggregation During Training: During model aggregation, FedGraph encrypts each client’s model updates before transmission. The central server then performs homomorphic aggregation on the encrypted updates to compute the global model without accessing any plaintext data (Figure 4(Zhang et al., 2020a; Jin et al., 2023).
Figure 5: Comparison of FedGCN performance on plaintext (blue) and with HE encryption (green).
Homomorphic Encryption Overhead: Figure 5 compares the training time and communication cost for the FedGCN method with and without homomorphic encryption. Since FedGCN involves both pre-training and training-phase aggregation, it serves as a representative case for evaluating the overhead introduced by HE across the full pipeline. As shown in the figure, HE significantly increases communication overhead, particularly during the pre-training phase, where feature matrices are often substantially larger than model parameters. These findings motivate our exploration of communication-efficient techniques in Section 4. Detailed benchmarks, including HE configurations and performance breakdowns, are in Appendix F.
@@ 3.3 FedGraph Kubernetes
Figure 6: Kubernetes Service for FedGraph Cluster Management.
To support diverse training workloads and scale to a large number of clients, FedGraph employs Kubernetes as the foundation of its infrastructure layer. Unlike prior platforms (Table 1) that lack native Kubernetes integration, FedGraph is designed to scale efficiently to both large client populations and massive graph datasets. As shown in Figure 6, FedGraph leverages Amazon Web Services Elastic Kubernetes Service to enable flexible and dynamic resource management. A self-managed Kubernetes cluster equipped with GPU instances is deployed to handle computationally intensive tasks. The architecture consists of a master node responsible for orchestration, task scheduling, resource allocation, and cluster health monitoring, while worker nodes execute distributed training processes in parallel. To enhance scalability and cost-efficiency, we integrate the Kubernetes Cluster Autoscaler, which dynamically adjusts worker nodes based on workload demands. This ensures resources are used only when needed, maintaining high efficiency under varying system loads for continuous large-scale execution.
3.4 Optimizations for Scalability
FedGraph is designed to support large-scale federated learning, handling graphs with thousands to millions of nodes through configurable training features.
Client Selection for Large-Scale Federated Learning: FedGraph uses a selective client engagement mechanism, allowing the user to specify a fraction of clients that send the model update at each training round (Li et al., 2019). This approach reduces communication costs at the server and resource usage at the clients, enabling scalability to thousands of clients.
Minibatch Training for Federated Updates: Each client processes only a subset of its local graph, reducing computation and memory demands. This also enables devices with limited resources relative to the size of their training datasets to participate in the training, improving convergence speed without overwhelming network or client resources.
Communication and Resource Optimization: FedGraph optimizes communication, minimizing data transfer between clients and the server. Kubernetes dynamically manages Server Pods and Trainer Pods to meet workload demands, ensuring efficient resource usage as the system scales.